OpenStudy (anonymous):

Identify the horizontal asymptote of f(x) = quantity 6 x minus 7 over quantity 11 x plus 8.

OpenStudy (anonymous):

a. 0 b. y=6/11 c. -7/8 d. no horizontal asymptote

OpenStudy (anonymous):

@dan815

OpenStudy (michele_laino):

I don't understand the formula for f(x), please can you write it better?

OpenStudy (anonymous):

6x - 7/11x + 8

OpenStudy (michele_laino):

if \[f(x)=6x-\frac{ 7 }{ 11x }+8\], we have: \[\lim _{x \rightarrow \pm \infty} f(x)=\pm \infty \] so we have not horizontal asymptotes

OpenStudy (anonymous):

okay thank you. very helpful

OpenStudy (michele_laino):

thank you!

OpenStudy (campbell_st):

there is a horizontal asymptote... a simple way to look at the problem the degree of the numerator and denominator are the same, the horizontal asymptote is the ratio of the coefficients \[y = \frac{6x - 7}{11x+ 8}\] both are degree 1 so the horizontal aysmptote is \[y = \frac{6}{11}\] to check graph it using https://www.desmos.com/calculator

OpenStudy (campbell_st):

I think the confusion has occurred because of the way the expressions have been written

OpenStudy (anonymous):

oh okay thanks